OpenStudy (anonymous):

Find the anti derivative of e^(-x^2)

OpenStudy (anonymous):

There isn't one "per se".

OpenStudy (anonymous):

How familiar are you with integration/differentiation? :P

OpenStudy (anonymous):

Well, i have learned differentiation already and so far i can find the indefinite integral and area under a curve. If that answers your question.

OpenStudy (turingtest):

If you are trying to show convergence or divergence for some interval, that can be done this integral cannot http://www.wolframalpha.com/input/?i=e%5E(-x%5E2)dx&t=crmtb01

OpenStudy (turingtest):

if it were perhaps xe^(-x^2) that would be very different

OpenStudy (anonymous):

Do you know about parametric differentiation under the integral sign turing?

OpenStudy (turingtest):

Leibniz rule stuff? a little bit only...

OpenStudy (anonymous):

i actually did go on that site before i went here. My main point in asking this question is because i have the find the area under two curves and y=e^(-x^2) is one of the equations. (The other one is y=x^3 -x +1)

OpenStudy (anonymous):

Well anyway, thank you Malevolence and TuringTest

OpenStudy (turingtest):

I think I have seen this kind of problem before. I have to think about that one though, pretty tricky I'll get back to you

OpenStudy (anonymous):

If you can that would be wonderful :) Would you like me to post it as a separate question instead?

OpenStudy (anonymous):

I was going to say, instead of \[\int\limits e^{-x^2}dx\] consider: \[\int\limits_{-\infty}^{\infty}e^{-\rho x^2}dx=\sqrt{\frac{\pi}{\rho}}, \rho >0. \] From this you can derive relations for: \[\int\limits_{-\infty}^{\infty}x^{2k}e^{- \rho x^2}dx \] for positive integers k. The odd powered ones go to zero (odd function over symmetric bounds).